OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (-3, -5) and (-11, -7)? (-8, -1) (-7, -6) (-4, -6)

OpenStudy (experimentx):

the second one

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

to find the mid point take the average in each coordinate

OpenStudy (anonymous):

\[\frac{-3-11}{2}=\frac{-14}{2}=-7\] is the first coordinate

OpenStudy (anonymous):

\[\frac{-5-7}{2}=\frac{-12}{2}=-6\]is the second coordinate